A Supermarket in Alabama occasionally receives a shipment of live shrimp from New Orleans.
Seventy-five percent of the time the store will place an ad and sell out of its shipment. The
store advertises the availability of shrimp for 82 percent of the shipments. What is the
probability that the store will sell an entire shipment if it advertises?
Answer #1

events :

A : advertises

S : sell out the shipment

P(A and S) = 0.75

P(A) = 0.83

Joint probability :

P(S and A) = P(S / A)*P(A)

0.75 = P(S / A)*0.83

P(S / A) = 0.75/0.83 = 0.9036 = 90.36%

probability that the store will sell an entire shipment if it advertises : P(S / A) = 90.36%

P.S. (please upvote if you find the answer satisfactory)